A wave goes through air with a velocity of 7.4 × 10^7ms-1. Find the velocity of the wave if the temperature of the air is reduced by a factor of 3


1
Expert's answer
2021-02-02T07:08:44-0500

Answer

Relationship between velocity frequency and temperature is given

"v \propto \nu \propto \frac{1}{T}"

"v_2=v_1 \frac{T_1}{T_2}=\frac{ 7.4 × 10^7 T_1}{3T_1}\\=2.46× 10^7 m/s"